OpenStudy (ttop0816):

please help!!!Solve sec^2θ + tanθ - 3 = 0 for 0° ≤ θ < 90°. θ = ???

OpenStudy (mertsj):

Replace sec^2 with 1 + tan^2

OpenStudy (mertsj):

\[1+\tan ^2\theta + \tan \theta -3=0\] \[\tan ^2\theta +\tan \theta -2=0\]

OpenStudy (mertsj):

Factor and set each factor equal to 0

OpenStudy (ttop0816):

would it beθ = 45 perhaps??

OpenStudy (mertsj):

I would think there would be more than 1 answer. Did you factor?

OpenStudy (mathmale):

ttop: Try factoring x^2 + x - 2 = 0 as an example of how to approach this problem. Then do the same thing to tan2θ+tanθ−2=0.